OpenStudy (anonymous):

What is the slope of the line passing through the points (0, –5) and (4, 2)? help

OpenStudy (cybershadow):

it cud be done easily , slope of the line is = (y2-y1)/(x2-x1)

OpenStudy (cybershadow):

u can put in the values and find it :)

OpenStudy (anonymous):

kk its 1/2?

OpenStudy (cybershadow):

y1 = -5,y2 = 2 ,x1 =0 ,x2=4 so slope = (2-(-5))/(4-0)= 7/4

OpenStudy (anonymous):

oh my bad What is the slope of the line whose equation is 2x – 4y = 10? thats the one

OpenStudy (cybershadow):

first of all we gotta convert it into the form y = mx +c

OpenStudy (cybershadow):

so rearranging the equation y = (2x-10)/4

OpenStudy (cybershadow):

that will make y = 1/2x -10/4

OpenStudy (cybershadow):

compare it with y = mx +c

OpenStudy (cybershadow):

it will give m = 1/2 which is the slope of the line :)

OpenStudy (cybershadow):

is it clear?

OpenStudy (anonymous):

yes tnakyou